Science Prep Night 20130208 UT

Summary

See: https://docs.google.com/a/lbto.org/document/d/1pMSQDB_Bx9xNWNOemAB_g10_Tt7O0hHxlJ3p5FG7cWU/edit

  • MODS - Restarted the ICs. Also needed to turn on tcs service and red caliban. Major delay was a problem with the laser power not turning on. See IT #4626. One primary spectrophotometric standard acquired and observed. Instrument seems to be in working order although IMCS and red caliban will need to be restarted tomorrow. Masks installed but maskSnap.pro was NOT executed. Olga will try to get to this tomorrow. I'll clean up the data when I get there.

  • LBC - Ran dofpia and Christian's data successfully. Non-sidereal targets, so further test of NSIGUI. Seems to all be behaving nicely.

  • LUCI - NA

MODS

02:02 UT -- Asked David to reboot ICs

02:07 UT -- Turned on MODS with modsWake. Tried a sieveSnap, but immediately had IMCS failure. We also saw that the data label was A://mods.20130430.00XX.fits

02:52 UT -- Tried to troubleshoot for quite a while in twilight. Turned instrument off and on. Tried to restart IMCS service. Tried to check laser power levels --> the set power was correct but the power was clearly off. Finally called Olga after exhausting my troubleshooting. Thought there might have been a problem with the IEB, but a reset proved to clear the error.

03:15 UT -- Olga arrives and we find a slew of little problems. TCS service was not running and neither was red caliban (although Olga saw this running earlier in the day). Note to everyone: computer = mods does NOT show data information like caliban etc. Must use mods1data for this

~~ 03:35 UT -- Still did not solve our problems. Able to see images, but no IMCS lock frown, sad smile

~~ 03:40 UT -- Reset of all MODS services

04:00 UT -- Olga digging through mails finds similar problem from last year (March 2012). Rick sent a solution, implemented it and it worked

04:15 UT -- Finally able to take images. Sievesnap, etc. all working correctly now

04:24 UT -- David collimates MODS

04:35 UT -- Start checkout of instrument and spectrophotometric data. MODS performing ok now. Checked out for science.

LBC

05:10 UT -- Instrument switch to LBC

05:24 UT -- Instrument switch complete

05:30 UT -- Turnover to Christian, dofpia, etc.

05:50 UT -- Science!! Christian starts with non-sidereal observing.

-- MichelleEdwards - 30 Apr 2013
Topic revision: r1 - 30 Apr 2013, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback